Two jets are traveling toward each other and are 4000km apart. The rate of one jet is 100km/hr faster than the rate of the other. If the jets pass each other after 2.5 hours, what is the rate of each jet?
Answer by stanbon(75887)

Two jets are traveling toward each other and are 4000km apart. The rate of one jet is 100km/hr faster than the rate of the other. If the jets pass each other after 2.5 hours, what is the rate of each jet?
---
closing speed for the two jets is x + x+100 = 2x+100 km/h
------------
Equation:
time = distance/rate = 4000/(2x+100) hrs = 2.5 hrs
---
2000/(x+50) = 2.5
2.5x + 125 = 2000
2.5x = 1875
---
x = 750 km/h (speed of the slower jet)
x+100 = 850 km/h (speed of the faster jet)
=============================================
Cheers,
Stan H.